OpenStudy (anonymous):

(2014k + 1)(2018k + 1) is a perfect square number. find the smallest value of k with k is natural number

OpenStudy (anonymous):

can anyone here help me in my ques

OpenStudy (thomas5267):

I don't see how this is possible unless 1 is a perfect square.

OpenStudy (thomas5267):

So k=0 I guess?

OpenStudy (rational):

(2014k + 1)(2018k + 1) = y^2 2014*2018k^2 + (2014+2018)k = y^2 - 1 4k(19*53*1009k + 3^2*7) = (y-1)(y+1)

OpenStudy (rational):

it is easy to see that "y" must be odd 0 is not natural number

OpenStudy (thomas5267):

So how do you solve for y?

OpenStudy (thomas5267):

Does this has anything to do with modular arithmetic?

OpenStudy (amorfide):

you can not say 0 is not a natural number as it is not universally agreed, some use 0 to be natural, some do not

OpenStudy (rational):

agree

OpenStudy (thomas5267):

Which grade is the OP in? Is this grade 12 mathematics or is it university number theory?

OpenStudy (rational):

he asks olympiad level problems always

OpenStudy (rational):

im getting k = 945

OpenStudy (raden):

if i resubtituting k = 945 to original equation. it does not works

OpenStudy (raden):

looks there is no k as the solution

OpenStudy (raden):

(2014k + 1)(2018k + 1) = y^2 2014*2018k^2 + (2014+2018)k + 1 = y^2 2014*2018k^2 + 4032k + 1 - y^2 = 0 D = x^2 b^2 - 4ac = x^2 (4032)^2 - 4(2014*2018)(1 - y^2) = x^2 16 + 4y^2 = x^2 x^2 - 4y^2 = 16 (x + 2y)(x - 2y) = 16 cases : x + 2y = 16 x - 2y = 1 or x + 2y = 8 x - 2y = 2 the 2 equations above cant be integer for y

OpenStudy (raden):

it means there is no k (natural number) as the solution which make (2014k+1)(2018k+1) is a perfect square number

OpenStudy (thomas5267):

x=2016 is such value. I brute-forced it lol.

OpenStudy (raden):

except x = 0 would be satisfied like you said @thomas5267 :D

OpenStudy (rational):

actually i worked a solution, i might be having some algebra error one sec..

OpenStudy (raden):

i checked it by using calculator if k = 945 (2014k + 1)(2018k + 1) = y^2 (2014(945) + 1)(2018(945) + 1) = y^2 y = 1905120.0625 bla bla bla :D

OpenStudy (thomas5267):

k=2016 is one answer.

OpenStudy (raden):

ah, yes k = 2016 is work.... my job is mistake also lol

OpenStudy (rational):

yeah algebra mistake, k=2016 is the only solution

OpenStudy (rational):

let me know if anyone is interested in the solution it is pretty long but i found it interesting when solving

OpenStudy (raden):

i think i found my mistake i just miss one extra case : (2014k + 1)(2018k + 1) = y^2 2014*2018k^2 + (2014+2018)k + 1 = y^2 2014*2018k^2 + 4032k + 1 - y^2 = 0 D = x^2 b^2 - 4ac = x^2 (4032)^2 - 4(2014*2018)(1 - y^2) = x^2 16 + 4y^2 = x^2 x^2 - 4y^2 = 16 (x + 2y)(x - 2y) = 16 cases : x + 2y = 16 x - 2y = 1 or x + 2y = 8 x - 2y = 2 or x + 2y = 4 x - 2y = 4 the third equation is work, get solution for (x,y) = (4,0) at least, subtitute y = 0 into : 2014*2018k^2 + (2014+2018)k + 1 = y^2 2014*2018k^2 + (2014+2018)k + 1 = 0^2 2014*2018k^2 + (2014+2018)k + 1 = 0 from here get k = 2016

OpenStudy (rational):

D = x^2 guarantees rational solutions, but never integral solutions right ?

OpenStudy (raden):

but if D not equal 0 be guarante the roots cant be integer :D

OpenStudy (raden):

i mean D not equal x^2

OpenStudy (rational):

Ahh I see, your method works !! XD

OpenStudy (rational):

pretty clever !

OpenStudy (raden):

i get this method from one of the member OS here, his name @mukushla my friend from iran

OpenStudy (rational):

he is my best friend on OS too

OpenStudy (raden):

really, we all are friend :D

OpenStudy (rational):

im not seeing him much these days looks he got busy after starting Msc...

OpenStudy (raden):

yeah, i didnt see if he was online on facebook